# Graph atomic cluster expansion for foundational machine learning interatomic potentials

### v1

@@ II.5.1 Al-Li
For fine-tuning, we utilized a dataset for the Al-Li binary system from Ref. Menon et al. (2024). We curated the data by removing structures from the liquid phase and those corresponding to randomly sampled space groups. This resulted in a total of three thousand structures within 1 eV/atom above the convex hull, from which we allocated 5% as a test set. To study data efficiency, we then created a series of training subsets of varying sizes (5%, 10%, 25%, 50%, and 75%) from the remaining data.
Using these training and test sets, we fine-tuned the 1L-OMAT and 2L-OMAT models by updating all weights, a process we term ”naive fine-tuning.” For comparison, we also trained two models from scratch: a one-layer ACE model as implemented in the PACE software Lysogorskiy et al. (2021); Bochkarev et al. (2022b) and a GRACE-2L model, designated 2L-baseline, with a complexity identical to GRACE-2L-OMAT.
@@ II.6 Model distillation
While foundational MLIPs are highly accurate and cover a wide range of chemical and configurational space, they often have lower computational performance and limited parallelization compared to chemistry-specific MLIPs. To overcome these limitations, model distillation can be used Morrow and Deringer (2022); Gardner et al. (2025). It involves retraining a simpler, faster ”student” model on a smaller, subsampled dataset that has been labeled by a foundational MLIP ”teacher” model.
To demonstrate this capability, we fine-tuned the GRACE-2L-OMAT model on the combined HEA25 Lopanitsyna et al. (2023b) and HEA25S Mazitov et al. (2024) datasets. We then performed distillation to the GRACE-FS model, an architecture chosen for its straightforward parallelization and CPU-only inference.
The initial foundational potential, 2L-OMAT, exhibits low formation energy MAE of 19.7 meV/atom for unary and 14 meV/atom for binary structures (MatProj, non-magnetic columns in Table 3). These numbers are comparable to the 23 meV/atom formation energy MAE on the Matbench Discovery leaderboard. Structures in the MatProj validation set with magnetic elements (Fe, Co, Ni, and Cr) were not considered due to the different settings for magnetism in MatProj (spin-polarized, PBE) and HEA25S (non-magnetic, PBEsol). However, even for non-magnetic elements, the different DFT functionals resulted in large prediction errors.
After fine-tuning, the resulting 2L-OMAT-HEA25S-ft model demonstrates a small MAE of 3.5 meV/at for energy and 85 meV/Åfor forces on the HEA25S validation set. However, its formation energy errors on the MatProj validation set increased, from 19.7 to 23 meV/atom for unaries and 14 to 30 meV/atom for binaries, due to the different DFT functionals.
This fine-tuned model was then used to label a synthetic training dataset, which comprised structures from HEA25 and HEA25S, along with unary and binary structures (both ideal and rattled configurations) for all 25 elements. Subsequently, the GRACE-FS model was parameterized using this synthetic data, resulting in the FS-distill model. For comparison, a GRACE-FS model of identical complexity, termed FS-baseline, was trained directly on the combined HEA25 and HEA25S datasets with their original DFT labels.
The FS-distill model exhibits poorer metrics on the HEA25S validation set (9.9 meV/at for energy and 158 meV/Åfor forces) compared to 2L-OMAT-HEA25S-ft, primarily due to its simpler ”student” model architecture. The corresponding error metrics on HEA25S for the FS-baseline model (8.8 meV/at for energy and 149 meV/Åfor forces) are slightly better than those of FS-distill. Both FS models demonstrate similar energy MAE to the models from the original Ref. Mazitov et al. (2024).
Despite these observations, the FS-distill model demonstrates notable advantages. It is not only able to capture increased chemical complexity, as indicated by its validation metrics on HEA25S, but also achieves significantly better unary and binary formation energy MAE (28.5 meV/atom and 33.6 meV/atom, respectively) compared to the FS-baseline model (36.2 meV/atom and 54.4 meV/atom), which was trained solely on the HEA25+HEA25S dataset.
Thus, using the distillation approach, that we combined here with fine-tuning the ”teacher” model, allows for obtaining specialized models with much simpler architectures and higher computational performance.
Table 3: Distillation of GRACE models. Error metrics for: (1) Formation energies (meV/atom) of unary and binary structures for Materials Project (PBE, spin-polarized) on 19 non-magnetic elements. (2) Energy(meV/at) and forces (meV/Å) for HEA25S (PBEsol, non-magnetic), 25 elements. See text for details.
Model MatProj, non-magnetic HEA25S validation Training set
unary binary
$E_{f}$-MAE E-MAE F-MAE
2L-OMAT 19.7 14 - - OMat24
2L-OMAT-HEA25S-ft 23 30 3.5 85 HEA25+HEA25S
FS-distill 28.5 33.6 9.9 158 Synth. HEA25+HEA25S+Unary/Binary
FS-baseline 36.2 54.4 8.8 149 HEA25+HEA25S
@@ IV.1 Models architecture and parameterization
All GRACE models were implemented in the grace-tensorpotential package, which is based on the TensorFlow library Abadi et al. (2016). The models implement a graph extension of the atomic cluster expansion (ACE) method, utilizing both star- and tree-like many-body basis functions, that form an orthonormal and complete basis set Bochkarev et al. (2024). The configurations for the GRACE foundation potentials together with parametrization details are outlined in Supplementary Information IX.1.
@@ IV.2 Validation
Foundational machine learning interatomic potentials. For the MACE models, the ’medium-mpa-0’ checkpoint was used for MACE-MPA-0, and ’mace-omat-0-medium’ for MACE-OMAT-0. All MACE models were using float64 precision and had the cuEQ option enabled. The MatterSim model utilized the ’mattersim-v1.0.0-5m’ checkpoint. The ORB model was based on the ’v3-conservative-inf’ modification, configured for ’float32-highest’ precision, that corresponds to full float32 precision. SevenNet employed the ’7net-mf-ompa’ checkpoint with its ’mpa’ modality. For eSEN, the ’esen_30m_omat’ checkpoint with a seed of 0 was selected. UMA-M and UMA-S models corresponded to their ’uma-m-1p1’ and ’uma-s-1p1’ checkpoints, respectively, specifying the ’omat’ task. The DPA model used the ’2025-01-10-dpa3-openlam’ checkpoint, whereas for computational performance tests “DPA-3.1-3M-ft” checkpoint was used. For eqV2_M model “eqV2_86M_omat_mp_salex” checkpoint was used. All GRACE models consistently operated with float64 precision.
@@
Elastic constants. We used elastic properties from the Materials Project De Jong et al. (2015) as our reference, employing an energy-based method Golesorkhtabar et al. (2013) to compute the elastic tensor $C_{ij}$ in Voigt notation. Out of 10073 reference elastic matrix calculations, we could not compute elastic constants for all structures due to relaxation and convergence issues with some potentials. Therefore, we relied on a common subset of 7962 structures for which elastic tensors were successfully computed by all potentials. This subset serves as a robust and representative test set for evaluating elastic matrix predictions. The eSEN model was not validated due to its high computational expenses.
Grain boundaries. Grain boundary structures and their corresponding reference structures were sourced from the Crystalium project Zheng et al. (2020), a dataset closely related to the Materials Project Jain et al. (2013). The reference structures were fully relaxed using the BFGS method with a FrechetCellFilter, applying a relaxation criterion of 0.001 eV/Å. Grain boundary relaxation was carried out using the FIRE minimization algorithm, also with a FrechetCellFilter, enforcing a maximum force threshold of 0.01 eV/Å  and a limit of 500 optimization steps. For eqV2 model, relaxation criteria was loosen to 0.02 eV/Å  due to numerical instabilities. In total, 327 grain boundaries were initially computed across 58 different elements. To ensure consistency and avoid discrepancies, only structures with a grain boundary plane orthogonal to the z-direction were selected, resulting in a final set of 297 grain boundary structures.
Surfaces. Surface structures and their corresponding reference data were sourced from the Crystalium project Tran et al. (2016), which is associated with the Materials Project Jain et al. (2013). The same relaxation settings as for grain boundaries were applied. From an initial total of 1124 surface structures, only those with a surface plane orthogonal to the z-direction were selected, yielding a final set of 716 surface structures for analysis.
Self-interstitials and vacancies. Reference data for self-interstitial and vacancy formation energies were taken from Ref. Ma and Dudarev (2019) for 13 BCC metals and Ref. Ma and Dudarev (2021) for 13 FCC metals. All these reference values were computed using the PBE functional.
@@ VI Code availability
VII Acknowledgements
The authors acknowledge funding from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through SFB 1394 (Project No. 409476157). The authors gratefully acknowledge the computing time made available to them on the high-performance computer Noctua-1 and Noctua-2 at the NHR Center Paderborn Center for Parallel Computing (PC2). This center is jointly supported by the Federal Ministry of Research, Technology and Space and the state governments participating in the National High-Performance Computing (NHR) joint funding program (www.nhr-verein.de/en/our-partners). The authors also acknowledge the computing time made available on the Elysium HPC cluster at Ruhr-University Bochum.
VIII Author contributions
Conceptualisation and Project Administration: All authors. Y.L. and A.B. developed the software and parameterized the models. Writing—original draft: Y.L. Writing—review and editing: All authors. Resources and funding acquisition: Y.L., R.D.
IX Competing interests
@@
IX.1 Models architecture and parameterization
Configurations of GRACE foundational potential are given in Table S1. We used spherical harmonics up to $l_{\mathrm{max}}=4$. Eighty-nine chemical elements from the OMat, sAlex, and MPTraj datasets were embedded into a 128-dimensional space. Each model has a 6 Å  cutoff radius and uses Chebyshev polynomials as its radial basis. These polynomials are transformed into radial functions using a multi-layer perceptron (MLP) with two hidden layers, each containing 64 units. Basis functions up to the fourth product order are constructed recursively using sparse coupling operations. Their maximum angular momentum, $L_{\mathrm{max}}$, varies depending on the product order. For two-layer models, equivariant functions from all product orders are linearly combined to create “messages” or indicators Bochkarev et al. (2022a). These are then passed to the next layer, where the same recursive construction is performed. Finally, all invariant basis functions from all product orders of the first and second layers are linearly combined to generate a set of atomic densities. The atomic energy is computed as the sum of the first atomic density and the result of the remaining densities passed through another MLP. The total energy of the structure is the sum of all atomic energies. Forces are computed via the gradient of the total energy with respect to the atomic positions. All calculations are done in double precision (FP64). The models is just-in-time (JIT) compiled using XLA to achieve optimal performance. Further details on the GRACE models will be published in a separate work.
@@ IX.1 Models architecture and parameterization
The GRACE models were parameterized using the recently published OMat24 Barroso-Luque et al. (2024b) dataset in combination with the sAlex Schmidt et al. (2023); Wang et al. (2023) dataset and the MPTraj dataset (v2022.10.28) Jain et al. (2013); Deng et al. (2023b). OMat24 includes a wide range of structures, whereas sAlex and MPTraj include only relaxation trajectories. Raw VASP energies, forces and stresses were used for parameterizations.

### v2

@@ II.5.1 Al-Li
For fine-tuning, we utilized a dataset for the Al-Li binary system from Ref. menon2024electrons. We curated the data by removing structures from the liquid phase and those corresponding to randomly sampled space groups. This resulted in a total of three thousand structures within 1 eV/atom above the convex hull, from which we allocated 5% as a test set. To study data efficiency, we then created a series of training subsets of varying sizes (5%, 10%, 25%, 50%, and 75%) from the remaining data.
Using these training and test sets, we fine-tuned the 1L-OMAT and 2L-OMAT models by updating all weights, a process we term ”naive fine-tuning.” For comparison, we also trained two models from scratch: a one-layer ACE model as implemented in the PACE software Lysogorskiy21; Bochkarev22 and a GRACE-2L model, designated 2L-baseline, with a complexity identical to GRACE-2L-OMAT.
@@ II.6 Model distillation
While foundational MLIPs offer high accuracy across a vast chemical and configurational space, they often have lower computational performance and limited parallelization compared to chemistry-specific models. To bridge this gap, model distillation can be employed morrow2022indirect; gardner2025distillation, a process involving the retraining of a simpler, faster “student” model on a dataset labeled by a foundational “teacher” model. We investigate different distillation and fine-tuning pathways using the combined HEA25 lopanitsyna2023modeling and HEA25S mazitov2024surface datasets as a case study. We employ the foundational GRACE-2L-OMAT model as the “teacher” and the GRACE-FS architecture as the “student”, selected for its straightforward parallelization and CPU-only inference capabilities.
We evaluated performance based on two tasks with corresponding accuracy metrics. The primary task reflects model accuracy on the new downstream application, measured by the force component MAE on the HEA25S validation set. The secondary task serves as a proxy for the model’s retention of general chemical knowledge and stability, as the HEA25S dataset lacks pure unary or binary structures. This was measured by the MAE of formation energy of binaries compounds from the Materials Project comprised of non-magnetic elements covered in the HEA25 dataset.
The different pathways for the downstream task are illustrated in Fig. 10. The initial foundation model (“Foundation”) exhibits a large error on the primary task due to the DFT functional mismatch between its training data (PBE) and the HEA25S target (PBEsol). However, it retains a low error (14 meV/atom) on the secondary task—comparable to the 23 meV/atom formation energy MAE on the Matbench Discovery leaderboard.
We explored three distinct approaches to address this. In the first method, we fine-tuned the foundation model on the HEA25S dataset to create a new teacher (“Finetuned”). This teacher, having achieved the lowest error on the primary task, was then used to re-label the HEA25S dataset for parameterizing the GRACE-FS student model called “Naive Distilled”. A second pathway is opposite to the first: we re-labeled the HEA25S dataset using the original foundation model (GRACE-2L-OMAT) to parameterize a “Raw distilled” student model, which was subsequently fine-tuned using the original HEA25S dataset to produce the “Distilled/Finetuned” model. Finally, the third approach involved parameterizing the GRACE-FS model from scratch using the HEA25S dataset directly to create the “Bespoke” model. As shown in Fig. 10, all three approaches yield models with similar error metrics, though the bespoke model performs slightly better on the primary task due to the usage of the original DFT data.
To improve performance on the secondary task, while preserving good accuracy of the primary task, we generated a synthetic training dataset comprising structures from HEA25 and HEA25S, along with unary and binary structures (including both ideal and rattled configurations) for all 25 elements. This extended dataset was labeled by the GRACE-2L teacher (“Finetuned”) and used to parameterize a new student model (“Extended Distilled”). This model demonstrates significantly improved performance on the secondary task, close to the teacher model, while retaining metrics on the primary task that are only slightly worse to the bespoke model.
[…]
Figure 10: Comparison of distillation and fine-tuning strategies. The axes illustrate the trade-off between accuracy on the primary task (HEA25S force MAE) and general chemical stability (secondary task: formation energy MAE of non-magnetic binaries from the Materials Project). Computational performance for both GRACE-2L and GRACE-FS architectures was evaluated on a 10-core CPU and normalized per core. See text for details.
@@ IV.1 Models architecture and parameterization
All GRACE models were implemented in the grace-tensorpotential package, which is based on the TensorFlow library tensorflow. The models implement a graph extension of the atomic cluster expansion (ACE) method, utilizing both star- and tree-like many-body basis functions, that form an orthonormal and complete basis set bochkarev2024. The configurations for the GRACE foundation potentials together with parametrization details are outlined in Supplementary Information X.1.
@@ IV.2 Validation
Foundational machine learning interatomic potentials. For the MACE models, the ’medium-mpa-0’ checkpoint was used for MACE-MPA-0, and ’mace-omat-0-medium’ for MACE-OMAT-0. cuequivariance-ops-torch-cu12 0.3.0 cuequivariance-torch 0.3.0 All MACE models were using float64 precision and had the cuEQ option enabled. The MatterSim model utilized the ’mattersim-v1.0.0-5m’ checkpoint. The ORB model was based on the ’v3-conservative-inf’ modification, configured for ’float32-highest’ precision, that corresponds to full float32 precision. SevenNet employed the ’7net-mf-ompa’ checkpoint with its ’mpa’ modality. For eSEN, the ’esen_30m_omat’ checkpoint with a seed of 0 was selected. UMA-M and UMA-S models corresponded to their ’uma-m-1p1’ and ’uma-s-1p1’ checkpoints, respectively, specifying the ’omat’ task. The DPA model used the ’2025-01-10-dpa3-openlam’ checkpoint, whereas for computational performance tests “DPA-3.1-3M-ft” checkpoint was used. For eqV2_M model “eqV2_86M_omat_mp_salex” checkpoint was used. All GRACE models consistently operated with float64 precision.
@@
Elastic constants. We used elastic properties from the Materials Project de2015charting as our reference, employing an energy-based method golesorkhtabar2013elastic to compute the elastic tensor $C_{ij}$ in Voigt notation. Out of 10073 reference elastic matrix calculations, we could not compute elastic constants for all structures due to relaxation and convergence issues with some potentials. Therefore, we relied on a common subset of 7962 structures for which elastic tensors were successfully computed by all potentials. This subset serves as a robust and representative test set for evaluating elastic matrix predictions. The eSEN model was not validated due to its high computational expenses.
Grain boundaries. Grain boundary structures and their corresponding reference structures were sourced from the Crystalium project zheng2020grain, a dataset closely related to the Materials Project jain2013commentary. The reference structures were fully relaxed using the BFGS method with a FrechetCellFilter, applying a relaxation criterion of 0.001 eV/Å. Grain boundary relaxation was carried out using the FIRE minimization algorithm, also with a FrechetCellFilter, enforcing a maximum force threshold of 0.01 eV/Å  and a limit of 500 optimization steps. For eqV2 model, relaxation criteria was loosen to 0.02 eV/Å  due to numerical instabilities. In total, 327 grain boundaries were initially computed across 58 different elements. To ensure consistency and avoid discrepancies, only structures with a grain boundary plane orthogonal to the z-direction were selected, resulting in a final set of 297 grain boundary structures.
Surfaces. Surface structures and their corresponding reference data were sourced from the Crystalium project tran2016surface, which is associated with the Materials Project jain2013commentary. The same relaxation settings as for grain boundaries were applied. From an initial total of 1124 surface structures, only those with a surface plane orthogonal to the z-direction were selected, yielding a final set of 716 surface structures for analysis.
Self-interstitials and vacancies. Reference data for self-interstitial and vacancy formation energies were taken from Ref. ma2019universality for 13 BCC metals and Ref. ma2021nonuniversal for 13 FCC metals. All these reference values were computed using the PBE functional.
@@ VI Code availability
VII Funding
We acknowledge funding from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through SFB 1394 (project number 409476157).
VIII Acknowledgements
High-performance computing (HPC) resources were provided by the Paderborn Center for Parallel Computing (PC2) and the Elysium HPC cluster at Ruhr-University Bochum.
IX Author contributions
Conceptualisation and Project Administration: All authors. Y.L. and A.B. developed the software and parameterized the models. Writing - original draft: Y.L. Writing-review and editing: All authors. Resources and funding acquisition: Y.L., R.D.
X Competing interests
@@
X.1 Models architecture and parameterization
The configurations of the GRACE foundational potentials are schematically depicted in Fig. S1 and provided in more detail in Table S1.
The atomic structure is represented by chemical species types $\mu_{i}$ and atomic positions $\mathbf{r}_{i}$, which are transformed into bond vectors $\mathbf{r}_{\mathrm{ij}}$ between neighboring atoms within a 6 Å  cutoff radius. Geometric information is encoded using a Chebyshev radial basis and spherical harmonics $Y_{\mathrm{lm}}$ up to $l_{\mathrm{max}}=4$ Drautz19; Lysogorskiy21; Bochkarev22. The Chebyshev polynomials are transformed into radial functions $R_{\mathrm{nl}}^{(\ldots)}$ using a multi-layer perceptron (MLP) with two hidden layers, each containing 64 units. Chemical species are embedded into a 128-dimensional space $Z_{i}$. Single-particle basis functions are constructed from these inputs and summed into atomic bases $A_{\mathrm{i,nl}}$ Drautz19; bochkarev2024. Many-body basis functions up to the fourth product order are constructed via recursive ACE evaluation using sparse coupling operations (see Fig. S1, bottom panel). During the recursive ACE basis evaluation, we utilize an equivariant sum operation $\oplus$, summing equivariant quantities with the same $l$-character, and a sparse equivariant coupling $\otimes$ operation which employs Clebsch-Gordan coefficients. We employ a specific coupling order to avoid degenerate product functions. The maximum angular momentum, $L_{\mathrm{max}}$, varies depending on the product order. In the one-layer model, this procedure yields a star-graph ACE basis. For two-layer models, equivariant basis functions are linearly combined to define equivariant atomic representations $I_{\mathrm{i,nL}}$ bochkarev2022multilayer. These representations carry geometric information about the atomic environment, effectively extending the interaction range to $2r_{\mathrm{cut}}$. They serve as input for a second recursive ACE evaluation, resulting in a tree graph GRACE basis bochkarev2024. Finally, all invariant basis functions from all product orders of the first and second layers are aggregated to form the complete GRACE basis, which is then linearly combined to generate a set of atomic densities $\varphi_{i,p}$ . The atomic energy is computed via an embedding scheme as the sum of the first atomic density and the result of processing the remaining densities through another MLP. The total energy of the structure is the sum of all atomic energies. Forces are computed via the gradient of the total energy with respect to the atomic positions. All calculations are performed in double precision (FP64). The models are just-in-time (JIT) compiled using XLA to achieve optimal performance. Further details on the GRACE models will be published in a separate work.
Figure S1: GRACE models: (top panel) overall scheme and (bottom panel) recursive ACE basis evaluation. See text for more details.
@@ X.1 Models architecture and parameterization
The GRACE models were parameterized using the recently published OMat24 barroso_omat24 dataset in combination with the sAlex schmidt2023machine; wang2023symmetry dataset and the MPTraj dataset (v2022.10.28) jain2013commentary; deng2023chgnet. OMat24 includes a wide range of structures, whereas sAlex and MPTraj include only relaxation trajectories. Raw VASP energies, forces and stresses were used for parameterizations.
@@
X.2 Model Distillation
Multiple fine-tuning and distillation pathways were evaluated, as detailed in Table S2 and illustrated in Figure S2.
Computational performance benchmarks were conducted on an Intel(R) Xeon(R) W-2255 CPU (3.70 GHz) equipped with 10 physical cores. We measured the wall-clock run time within LAMMPS for a system containing 2000 tungsten atoms in a body-centered cubic (BCC) configuration. For GRACE-2L models, LAMMPS MPI parallelization was disabled; instead, we relied on TensorFlow’s internal parallelization with XLA compilation to utilize the available cores. The environment variables were configured as follows: OMP_NUM_THREADS=1, TF_NUM_INTRAOP_THREADS=20, and TF_NUM_INTEROP_THREADS=1. For GRACE-1L models, a hybrid strategy was employed to saturate CPU capacity, combining LAMMPS MPI parallelization (10 MPI ranks) with XLA/TensorFlow internal parallelization, utilizing TF_NUM_INTRAOP_THREADS=8 and TF_NUM_INTEROP_THREADS=1. Finally, for GRACE-FS, parallelization was achieved exclusively using 10 MPI ranks in LAMMPS. For comparison, we also measured computational performance on a single NVIDIA A100 80Gb GPU. GRACE-FS models of three different complexities were considered: large, medium, and small. We initiated our study with both GRACE-2L-OMAT and GRACE-1L-OMAT foundation models. These models were fine-tuned on the HEA25S dataset, resulting in the 2L-OMAT-HEA25S-ft and 1L-OMAT-HEA25S-ft models. As expected, the 2L model demonstrated better metrics after fine-tuning compared to the 1L model. These models were then distilled into two corresponding GRACE-FS-large models (FS-large-distill-1L/2L-OMAT-ft-HEA25s), implementing a naive distillation scenario. Interestingly, the accuracy of the distilled models does not depend on the distillation source, suggesting that the student models have reached their learning capacities. We also parameterized three bespoke GRACE-FS models of varying complexities. As expected, the accuracy of these models increases with model complexity. Additionally, we evaluated another sequence: first distilling the foundation models into student GRACE-FS models using the HEA25S dataset, followed by fine-tuning the students using the original HEA25S dataset with DFT references. As shown in the results, all three pathways yielded models of similar accuracy in both primary and secondary tasks, although the latter performed slightly worse compared to the foundation and fine-tuned models. To improve the performance of distilled models on the secondary task, we extended the distillation dataset with unary and binary compound structures (both ideal and rattled). This allowed us to parameterize GRACE-FS models across three different complexities, reaching secondary task accuracy levels comparable to fine-tuned models while only slightly compromising accuracy on the primary task. Increasing the complexity of both student and teacher models systematically improved metrics on both primary and secondary tasks (Fig. S2, left panel). The Pareto front of speed-accuracy, as shown in Fig. S2 (middle and right panels), is formed by the fine-tuned foundation 1L- and 2L-models, as well as the distilled models.
Finally, we performed an extended distillation where, in addition to total energies, forces, and stresses, atomic site energies were included as reference targets (FS-large-extdistill-atomic-energy-2L-OMAT-ft-HEA25s). However, the resulting metrics were slightly worse than those obtained from the identical parameterization without atomic energies (FS-large-extdistill-2L-OMAT-ft-HEA25s).
[…]
Model Training set Reference method HEA25S MatProj $t_{\mathrm{LAMMPS}}$
E-MAE F-MAE unary binary A100 10xCPU
Foundation
2L-OMAT OMat24 PBE 617.9 193.5 19.7 14 25 3448
1L-OMAT OMat24 PBE 615 207.1 23.9 15.6 10 494
Finetuned
2L-OMAT-HEA25S-ft HEA25S PBEsol 3.5 85 23 30 25 3448
1L-OMAT-HEA25S-ft HEA25S PBEsol 4.1 102.6 28.0 34.6 10 494
Naive distilled
FS-large-distill-2L-OMAT-ft-HEA25s HEA25S 2L-OMAT-ft-HEA25s 7.5 140.0 34.1 53.0 8 61
FS-large-distill-1L-OMAT-ft-HEA25s HEA25S 1L-OMAT-ft-HEA25s 7.5 140.1 34.1 52.2 8 61
Bespoke
FS-large-HEA25s HEA25S PBEsol 7.6 140.7 30.71 48.56 8 61
FS-medium-HEA25s HEA25S PBEsol 9.4 149.7 34.9 56.0 8 46
FS-small-HEA25s HEA25S PBEsol 9.9 157.6 35.7 55.6 8 46
Distilled/Finetuned
FS-large-ft-HEA-2L-OMAT HEA25S 2L-OMAT - 219.9 32.8 43.8 8 61
FS-large-ft-HEA-2L-OMAT-ft-HEA25s HEA25S PBEsol 8.2 144.5 34.1 51.9 8 61
FS-large-ft-HEA-1L-OMAT HEA25S 1L-OMAT - 224.2 31.4 39.9 8 61
FS-large-ft-HEA-1L-OMAT-ft-HEA25s HEA25S PBEsol 8.3 145.0 33.1 49.1 8 61
Extended distilled
FS-large-extdistill-2L-OMAT-ft-HEA25s ext. HEA25S 2L-OMAT-ft-HEA25s 9.9 158 28.5 33.6 8 61
FS-large-extdistill-1L-OMAT-ft-HEA25s ext. HEA25S 1L-OMAT-ft-HEA25s 10.1 154.4 30.3 36.7 8 61
FS-medium-extdistill-2L-OMAT-ft-HEA25s ext. HEA25S 2L-OMAT-ft-HEA25s 11.5 165.8 30.7 34.8 8 46
FS-medium-extdistill-1L-OMAT-ft-HEA25s ext. HEA25S 1L-OMAT-ft-HEA25s 11.6 163.9 30.2 38.5 8 46
FS-small-extdistill-2L-OMAT-ft-HEA25s ext. HEA25S 2L-OMAT-ft-HEA25s 12.0 170.8 30.8 34.5 8 46
FS-small-extdistill-1L-OMAT-ft-HEA25s ext. HEA25S 1L-OMAT-ft-HEA25s 11.6 167.9 32.2 40.3 8 46
Extended distilled (atomic energies)
FS-large-extdistill-atomic-energy-2L-OMAT-ft-HEA25s ext. HEA25S (+atomic energies) 2L-OMAT-ft-HEA25s 10.9 159.6 29.3 34.1 8 61
Figure S2: Performance trade-offs for GRACE models trained via different strategies. Left: Accuracy on the primary task (MAE of forces for HEA25S validation set) versus accuracy on the secondary task (MAE of formation energy of unaries and binaries from Materials Project). Lower values on both axes indicate better performance. Middle and Right: Computational cost (wall-clock time per atom step) measured on a single Intel(R) Xeon(R) W-2255 CPU with 10 physical cores, as normalized per core versus primary (middle) and secondary (right) tasks accuracies. The Pareto front (dashed line) highlights the optimal trade-off between speed and accuracy. Symbols differentiate model architectures (squares/diamond: GRACE-1L/2L models; and GRACE-FS models: circles - large, up triangles - medium, down triangles - small), while colors indicate the training methodology. Empty symbols indicate distilled models with 1L teacher, whereas filled symbols - 2L teacher.
X.3 Molecular dynamics stability and energy conservation
Figure S3: Total energy conservation for a 3,000-atom FLiBe supercell simulated in the NVE ensemble at 973 K using the GRACE-2L-OMAT-L model. The main plot shows the deviation of the total energy per atom from the mean over a 1 ns trajectory. The red line indicates the linear regression of the energy drift, calculated as $5\cdot 10^{-9}$ eV/atom/ns. The side panel shows the corresponding histogram of energy fluctuations, demonstrating a stable equilibrium distribution.
To assess the long-term numerical stability of the foundational models, we analyzed the conservation of total energy during a 1 ns molecular dynamics simulation of molten FLiBe in the NVE ensemble at 973 K. Figure illustrates the time evolution of the total energy fluctuations per atom. The system exhibits exceptional stability, characterized by a negligible linear drift of $5\cdot 10^{-9}$ eV/atom/ns. Furthermore, the histogram of energy fluctuations (right panel) displays a symmetric, Gaussian - like distribution centered at zero, confirming the absence of systematic heating or cooling artifacts over the simulation trajectory.
X.4 Detailed Validation Benchmarks
This section provides a detailed overview of the validation benchmarks.
Mean absolute error of formation energy from Matbench Discovery vs. computational performance are shown on Fig. S4. The performance for predicting the elastic tensor is shown in Fig. S5. For structural defects, error metrics for unary grain boundary formation energies are presented in Fig. S6, with a detailed element-wise decomposition of the SRME provided in Fig. S7. Similarly, the performance for unary surface formation energies is summarized in Fig. S8, with a corresponding elemental breakdown shown in Fig. S9. Finally, the model’s accuracy for point defects in unaries is presented for self-interstitials (Fig. S10) and vacancies (Fig. S11). The element-decomposed SRME for these point defects are detailed in Fig. S12 and Fig. S13, respectively.